Determine the quadratic function f whose graph is given.
The vertex is (4,-3) and the other point is (3,-1)
f(x)=a(x-h)^2+k
write answer in form: ax^2+bx+c
Answer= 2x^2-16x+29
Can someone break this down for me, especially how to get the 29.
Thank you.

Answer by ewatrrr(24785)


Hi
f(x)=a(x-h)^2+k V(4,-3)
f(x) = a(x-4)^2 -3 Using (3,-1) to solve for a
-1 = a - 3
2 = a
f(x) = 2(x-4)^2 -3
= 2(x^2 - 8x + 16) - 3
= 2x^2 - 16x + 32 - 3
f(x) = 2x^2 - 16x + 29